CENNET VARSA EDİSON GİDER

Beni Camiye çağırma hoca Fabrika olmalı, tütmeli baca Çalışmayı öner aça, muhtaca Avare kul ile bir etme beni

Haramdan bahseder yan gelir yatar Mevlide gider duayı satar Abdest, Namaz der de Cemaate çatar Takunya giydirip yürütme beni

İlim sahibine Gavurdur deme Azdırdın hoca sapıttın yine Milletin sırtına sen bine, bine Ben cahil değilim tor etme beni

Sakalı sıvazlar sahte şükreder Gider camide cennet vaad eder Cennet var ise Edison gider Beynimi yıkayıp delirtme beni

Edison cennete gidemez diyen Muhakkak yobazdır bu haltı yiyen Elinde tespihi, Cüppeyi giyen Yatırıp, kaldırıp çürütme beni

Edison şalteri indirdiği zaman Kesilir cereyan halimiz yaman Şaşırır kalırız işte o zaman Şamdanda mum diye eritme beni

Edison enerjim, gece güneşim İbadet sayarım var ise işim Benim meleğim evdeki eşim Sen yerin dibinde var etme beni

Düğmeye basıp da lamba yanınca O anda karanlık aydınlanınca Salavat getirir sarıklı hoca Gözümü bağlayıp kör etme beni

DOĞAN’ım sözlerim hayınlarına Aman siz gelmeyin oyunlarına Vebali, günahı boyunlarına Çağımızda yaya yürütme beni.

Halil Doğan, 20.09.1995 IF THERE’S A HEAVEN, EDISON IS GOING

Don’t call me to the prayer-house^{a}, Father^{b}.

It should be a factory; the chimney should smoke. For the hungry and the needy, propose work. Don’t take me for an idle wight.

He determines what is sinful as he lolls about.

He goes to funerals; he sells prayers.

He speaks of ablutions, of worship; he scolds the congregation. Don’t have me go about in clogs^{c}.

Don’t call the man of science an infidel.

Still you have corrupted, Father; you have led astray.

As you climb onto the backs of the people, I am not ignorant—don’t try to catch me.

He strokes his beard, he gives insincere thanks. In the prayer-house he goes and promises heaven. If there is a heaven, Edison is going. Don’t brainwash me, don’t make me go mad.

Whoever says Edison can’t go to heaven Is surely a fanatic—whoever utters this crap, Rosary in hand, on shoulders an alb. Don’t prostrate me and raise me, don’t putrefy me.

When Edison turns off the switch, The current is cut, we’re in dire straits, We’re bewildered at that moment. Don’t make me melt away like a candle.

Edison is my energy, my night-time sun. I’m at divine service if I have work. My angel is my spouse at home. Don’t give me life in the grave.

When he presses the button and lights the lamp,

Then—when the darkness is illuminated— Father in his mitre^{d} says the last rites. Don’t bind my eyes, don’t blind me.

I am FALCON: my words are for the deceitful. For goodness’ sake don’t fall into their trap. The sin, the crime is on their heads. In this day and age, let me not go on foot.

English interpretation by Ayşe Berkman and David Pierce

http://metu.edu.tr/~dpierce/

aliterally, mosque blit., hodja, that is, (religious) teacher, &c.

csc. for ablutions dlit., turban, which is one sense of the Greek word mitra; Herodotus uses this word to describe (1.195) the Babylonian and (7.90) the Cyprian head-dress